MODS Calibration Lamp Change Log

Fill out the table below every time you change a calibration lamp, regardless of whether it is an emergency replacement on-telescope or a planned preventative-maintenance replacment.

A lamp exchange action is not complete until the log below is updated.

Note that if you replace one of the Pen-Ray wavelength comparison lamps, you need to file an IssueTrak report with next-action the MODS support astronomer to purchase and stock its replacement on the mountain.

Also remember to complete (and if possible, close) any open IssueTrak reports that required a lamp exchange as next-action.

Calibration Lamp Change Log

Add your new entry at the BOTTOM of the table. One entry per lamp changed. Designations:
  • VFLAT = incandescent variable-intensity flat-field continuum lamp
  • QTHn = Quartz-Tungsten-Halogen continuum lamp (note n=1 or n=2, e.g., QTH1)
  • Hg(Ar) = Hg(Ar) Pen-Ray comparison lamp
  • Ne = Neon Pen-Ray comparison lamp
  • Ar = Argon Pen-Ray comparison lamp
  • Xe = Xenon Pen-Ray comparison lamp
  • Kr = Krypton Pen-Ray comparison lamp
If there are any notes (e.g., "lamp blackened", "pen-ray lamp pigtail cord twisted/broken"), please note it in the last column

Date Instrument Calib Lamp Initials Notes
2012 Aug 14 MODS1 VFLAT & QTH1 rwp/ts Planned summer maintenance replacement
2015 Sep 27 MODS2 Ne JP Burnt out? Removed source placed in box labeled POGGE - MODS TOOLS in cabinet #3 on 3U.
2016 Jan 5 MODS2 Hg(Ar), Ar, Xe, Kr JP Exchanged lamps for the flight lamps into MODS2 (Ne was exchanged in Sept). Lab lamps in cabinet on 3U, marked with date.
2017 Aug MODS1/2 VFLAT dpp Changed VFLAT lamps on MODS1 and MODS2, documented procedure for handover documentation
2018 Sep MODS2 Ne es Changed Ne on MODS2 as it was noted to be slow to light
2018 Oct MODS2 VFLAT es Replaced VFLAT lamp on MODS2. Ongoing problems with lamp intermittently failing (IT 7009). Problem continues with new lamp
2018 Nov MODS1 QTH1 es QTH1 on MODS1 had burnt out (IT 7461).
2020 Jun 2 MODS1 Ne dr/opk Replaced Ne on MODS1 (IT 8107). Took Ne lamp from LUCI cabinet, no spare found in MODS cabinet
2021 Apr 6 MODS2 Ne jm Replaced Ne on MODS2 (IT 7869). Took Ne lamp from MODS spares cabinets on 3U, one spare remaining. Old lamp placed in box and labeled, placed with other removed lamps in spares box.
2022 Mar 3 MODS1 Ne cb,js Chris Bolyard replaced the Ne lamp and its power supply
2022 Mar 3 MODS2 Ne dr,js Dan Rapoza replaced the Ne lamp and its power supply
-- RichardPogge - 14 Aug 2012
Topic revision: r13 - 03 Mar 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback